OpenStudy (anonymous):

Multiply: (x^3+y^5)(x^3-y^5)

OpenStudy (ajprincess):

This of the form (a-b)*(a+b). (a-b)*(a+b)=a^2-b^2 Here a=x^3 and b=y^5 Can u do nw @Helenshouse?

OpenStudy (anonymous):

(x^3-b^5)*(x^3+y^5) ???

OpenStudy (anonymous):

x9 - y25

OpenStudy (ajprincess):

(a-b)*(a+b)=a^2-b^2 (x^3-y^5)*(x^3+y^5)=(x^3)^2-(y^5)^2 Does that make sense

OpenStudy (anonymous):

Just multiply exponents?

OpenStudy (anonymous):

yes u just have to multiply the exponents but use the given identity [by ajprincess]

OpenStudy (anonymous):

Got it Thank you both.

OpenStudy (ajprincess):

yw:) Bt can u plz tell me vat u get as ur final answer?

OpenStudy (ajprincess):

@Ekaansh it wud be really better if u check the answer u got.

OpenStudy (anonymous):

What do you mean?

OpenStudy (ajprincess):

I mean x^9-y^25 is not the final answer

OpenStudy (anonymous):

x^6-y^10

OpenStudy (ajprincess):

ya that's right:)

OpenStudy (anonymous):

I'm just not getting the hang of this......Dividing polynomials is an issue for me as well.

OpenStudy (anonymous):

I can not tell when to add or subtract and when to reverse the signs.......

OpenStudy (anonymous):

This one for example......Why do we subtract.

OpenStudy (ajprincess):

I am nt exactly following ur question. If u hav something in the form (a-b)(a+b) then it can be written as difference of squares. that is a^2-b^2. That's a general form.  Following me?